OpenStudy (anonymous):

Fill in the questions and find the rule. What should the rule of the table be if the in is 100 and the out is blank Then the in is 9 and the out is 0.9 Then the in is 50 and the out is 5 Then the out is 1.5 but the in is blank . Then again the in is missing but the out is 0.5

OpenStudy (anonymous):

9/10=0.9 50/10-5 15/10=1.5 5/10=0.5

OpenStudy (anonymous):

So the rule is /10?